A company is expected to pay dividend of $6, $6, $6.50 over the next three years. A dividend of $7 is expected to be paid in the 4th year and in subsequent years it is anticipated that dividend will grow at 0.05%. If the required rate of return is 12%. Determine the value of the company's share. Will long term growth rate add value to the company if it is financed by retaining 30% of it's retained earnings.


1
Expert's answer
2021-03-18T16:44:12-0400

Value of share

First discount the dividends using rate of 12%

Growth rate is 0.05%

value

"D1=6\times \frac{1}{1.12^1}=5.357"


"D2=6\times \frac{1}{1.12^2}=4.783"


"D3=6\times \frac{1}{1.12^3}=4.627"


D4 use the following formula:


"D4= \frac {Do(1+g)}{Ke-g}\times \frac{1}{1.12^3}=\frac {7(1+0.0005)}{0.12-0.0005} \times \frac{1}{1.12^3}=41.72"


Value of share = D1+D2+D3+D4


Value=5.357+4.783+4.627+41.72 = $ 56.49


Yes, long term growth rate will add value because retained earnings are a cheap source of capital and readily available for use.